A real estate agent received a 5% commission on the sale of a home. If her commission was $43150 how much did the home sell for?
Answer by Maths68(1474)

Let
Home sold for = $x

5% of x = 43150
5/100 (x) = 43150
5x/100=43150
x/20=43150
x=43150*20
x=863000
Home Sold for $863000